# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ 4 Computationally Efficient Algorithm by Online Learning
Notably, a similar confidence set is achieved by Hwang and Oh (2023) by using the MLE defined in (1). In contrast, we obtain the same results by using an online estimator which only suffers $\mathcal{O}(1)$ computation and storage cost per round. Besides, we identify a technical issue in their analysis. Specifically, they bound the confidence set using the self-normalized concentration for vector-valued martingales in Lemma 15. However, it is crucial to recognize that the noise is not independent and satisfies $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ since the learner visits each stage $h$ exactly once per episode. Thus, the noise $\epsilon_{i,h}^{s^{\prime}}$ becomes deterministic and non-zero given the remaining noise $\epsilon_{i,h}^{s^{\prime\prime}}$ for $s^{\prime\prime}\neq s^{\prime}$. This contravenes the zero-mean sub-Gaussian condition in Lemma 15. We observe similar oversights also appear in the works on multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). To our knowledge, this issue has not been explicitly identified in previous studies. We note that this issue can be resolved by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al. (2024b) with only slight modifications in constant factors.
Algorithm 1 UCRL-MNL-OL
1:  Initialization: $\widehat{\Sigma}_{1,h}=\lambda I,\widehat{\theta}_{1,h}=\mathbf{0}$ for all $h\in[H]$.
2:  for $k=1,\ldots,K$ do
3:     Compute $\widehat{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (7).
4:     for $h=1,\ldots,H$ do
5:         Observe current state $s_{k,h}$, select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
6:         Set $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
@@
7:         Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$.
8:     end for
9:  end for
Efficient Optimistic Value Function Construction.  Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, the most direct way is to adopt the principle of “optimism in the face of uncertainty” and construct the optimistic value function as the maximum expected reward over the confidence set. Specifically, we define the optimistic value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
However, this construction is not efficient as it requires solving a maximization problem over the confidence set. Actually, the algorithm only requires the estimated action-value function to be an upper bound for the true value functions. Thus, we can use a closed-form confidence bound instead of computing the maximal value over the confidence set. To this end, we present the following lemma that enables us to construct the optimistic value function more efficiently.
@@
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}.$
Based on this lemma, we can replace the maximization problem in (6) with closed-form confidence bound and construct the optimistic value function $\widehat{Q}_{k,h}(s,a)$ as
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\widehat{\Sigma}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (7)
At state $s_{k,h}$, our algorithm chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$. Algorithm 1 presents the detailed procedure. We show that our algorithm achieves the following regret guarantee.
Theorem 1.
For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following regret guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}\kappa^{-1}dH^{2}\sqrt{K}\big{)}.$
Remark 4.
Our algorithm attains the same regret of Hwang and Oh (2023), yet for the more general inhomogeneous MDPs. Importantly, our algorithm only requires constant computational and storage costs, matching the computational efficiency of the linear cases (Jin et al., 2020; Zhou et al., 2021).
5 Statistically Improved Algorithm by Local Learning
In this section, we present an enhanced algorithm, UCRL-MNL-LL, that leverages local information to improve the statistical efficiency for MDPs with MNL function approximation.
While the UCRL-MNL-OL algorithm in Section 4 offers favorable computational and storage efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret. By Assumption 1, the quantity $\kappa$ may be exponentially small in the worst case, exhibiting an exponential dependence on the radius of the parameter set and linear in the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$. This creates a significant gap between MNL and linear function approximation.
Recently, the improved dependence on $\kappa$ has been achieved in the logistic bandit literature (Faury et al., 2020; Abeille et al., 2021; Périvier and Goyal, 2022) by the use of generalization of the Bernstein-like tail inequality (Faury et al., 2020) and the self-concordant-like property of the log loss (Bach, 2010). Thus, a natural question then arises: Can we improve the statistical efficiency of MNL function approximation for MDPs? We answer this question affirmatively by proposing an enhanced algorithm that reduces the dependence on $\kappa$ significantly through the use of local information. Though the achievement has been made in the logistic bandit, the extension to MDPs with MNL function approximation is non-trivial and new techniques specific to MDPs are required.
We first analyze where the dependence on $\kappa$ comes from. By the analysis of Theorem 1, we can see that the regret of the algorithm can be upper-bounded as follows,
$\operatorname*{Reg}(K)\leq 2H\widehat{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (8)
Here, the first term corresponds to the overestimation of the value function, and the second term represents the martingale difference sequence arising from the stochastic transition dynamics. The second term can be bounded using the Azuma-Hoeffding inequality, which is independent of $\kappa$. For the first term, by the construction of the confidence set in Lemma 1, $\widehat{\beta}_{k}$ represents the width of the confidence set, contributing to a $\kappa^{-1/2}$ dependence. The feature covariance matrix $\widehat{\Sigma}_{k,h}$ defined in (4) reflects the degree of exploration across different states, which also results in a $\kappa^{-1/2}$ dependence.
To mitigate these dependencies, we design a local learning algorithm that: (i) constructs a confidence set independent of $\kappa$; and (ii) builds a $\kappa$-independent feature covariance matrix that effectively captures the exploration degree across different states. We introduce the details in the following.
5.1 Improved Online Estimation
In Property 1, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain, as in (3). This quantity measures the exploration degree across different states and is used to update the parameter estimation and construct the confidence set. However, the bound is not tight, as the Hessian matrix can be significantly larger in certain regions, away from the global minimum. This observation motivates us to design a local learning algorithm that can adaptively leverage local information for improved guarantee.
Inspired by recent advances in multinomial logistic bandit (MLogB) (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (MNL) (Lee and Oh, 2024), we run an online mirror descent algorithm to estimate the parameter $\theta_{h}^{*}$. Differently from (4), we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound. Specifically, we estimate $\widetilde{\theta}_{k,h}$ as follows:
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (9)
where $\eta$ is the step size, $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d}$ and ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$. Note that both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally. Similar to the update in (4), the optimization problem in (9) can be efficiently solved using the following two-step update:
$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\widetilde{\mathcal{H}}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\mbox{and}\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$
This two-step update procedure incurs a computational cost of $\mathcal{O}(d^{3}U)$ and a storage cost of $\mathcal{O}(d^{2})$ per episode, both independent of the episode count $k$.
Based on this estimator, we can construct the $\kappa$-independent confidence set as follows.
Lemma 3.
For any $\delta\in(0,1)$, define the confidence set as
@@ 5.1 Improved Online Estimation
Zhang and Sugiyama (2023) studied the multiple-parameter MLogB model, where the unknown parameter is a matrix. Consequently, the confidence set in Theorem 3 of their work exhibits a polynomial dependence on the number of possible outcomes, which corresponds to the number of reachable states $U$ in our setting. This dependence is acceptable in the bandit setting while is not suitable for the MDP setting. In contrast, Lee and Oh (2024) focused on the single-parameter MNL model and revisited the self-concordant-like property, demonstrating that the log-loss of the single-parameter MNL model is $3\sqrt{2}$-self-concordant-like (Proposition B.1 in Lee and Oh (2024)). This property is crucial for the improved confidence set in Lemma 3 that is independent of $\kappa$ and $U$.
Algorithm 2 UCRL-MNL-LL
@@
3:     Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (11).
@@
7:         Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$.
8:         Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
@@
5.2 Improved Optimistic Value Function Construction
Based on the confidence set in Lemma 3, a natural choice for the optimistic value function construction is analogous to (7) and can be expressed as:
$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
However, though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (10), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence. To address this issue, we propose a new construction of the optimistic value function. Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from transition estimation errors.
Although the idea of using second-order Taylor expansion has been explored in bandits (Périvier and Goyal, 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024), fundamental differences arise in the MDP setting. Specifically, Périvier and Goyal (2022) studied the uniform revenue setting, where the reward is identical for all actions. Zhang and Sugiyama (2023) and Lee and Oh (2024) focused on the non-uniform setting, but the rewards for different actions are known a priori. However, in the MDP setting, the value function is state-dependent and unknown to the learner, leading to a more challenging problem. Moreover, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward. Due to these differences, the analyses used in the bandit setting cannot be directly applied to the MDP setting.
Lemma 4.
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}.$
where
$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (11)
In contrast to the value function specified in (10), where the term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is utilized, the refined value function introduced in (11) substitutes this term with $\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}$. This adjustment better preserves local information, offering a more precise and $\kappa$-independent estimation error bound.
At state $s_{k,h}$, our algorithm chooses action $a_{k,h}=\max\widetilde{Q}_{k,h}(s_{k,h},\cdot)$. The detailed algorithm is presented in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL in the following theorem.
@@ 5.2 Improved Optimistic Value Function Construction
For any $\delta\in(0,1)$, set $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
@@
The high-order term in Theorem 2 is now independent $\kappa$, significantly improving the statistical efficiency compared with Theorem 1. In comparison with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear cases (Zhou et al., 2021; He et al., 2023), the higher-order term in Theorem 2 only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.

### v3

@@ 4.3 Regret Guarantee
Focusing on the dominant term, our guarantee eliminates the problematic dependence on $\kappa^{-1}$, in stark contrast to the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ result of Hwang and Oh (2023). As noted in Claim 1, such undesirable dependence has polynomial scaling with the number of reachable states $U$, which can be as large as the entire state space $S$ in the worst case. This renders the guarantee for function approximation—designed for settings with large state and action spaces—essentially vacuous.
5 Computationally Efficient Algorithm
While the UCRL-MNL-LL algorithm is the first statistically efficient algorithm for MNL mixture MDPs, it is computationally expensive dur the the optimization of the MLE in Eq. (1) and non-convex optimization in Eq. (6). To address these challenges, we propose a computationally efficient algorithm in this section, which attains the same regret but with constant computational costs per episode.
5.1 Efficient Online Parameter Estimation
In this section, we focus on estimating the unknown parameter $\theta_{h}^{*}$ in a computationally efficient manner. We first discuss the storage and time complexities of the MLE optimization in Eq. (1). Next, we introduce an efficient online parameter estimation based on online mirror descent that provides similar guarantees to the MLE, but with constant storage and time complexity per episode.
For the storage complexity, the optimization problem defined in Eq. (1) requires storing all historical data, resulting in a storage complexity of $\mathcal{O}(k)$ at episode $k$. In terms of time complexity, the problem does not have a closed-form solution and can only be solved to within an $\varepsilon$-accuracy, such as using projected gradient descent. As discussed in Faury et al. (2022), optimizing the MLE typically requires $\mathcal{O}(\log(1/\varepsilon))$ iterations to achieve an $\varepsilon$-accurate solution. Since the loss function is defined over all historical data, each gradient step incurs a query complexity of $\mathcal{O}(k)$. As $\varepsilon$ is usually chosen as $1/k$ for episode $k$, the total time complexity is $\mathcal{O}(k\log k)$ at episode $k$. Consequently, both storage and time complexities scale linearly with the episode count, which is computationally expensive.
To improve the computational efficiency, the basic idea is to estimate the unknown parameter with the online mirror descent (OMD) update instead of the MLE as defined in Eq. (1). To this end, we first define per-episode loss function $\ell_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
$\ell_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\nabla\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$ (7)
$H_{k,h}(\theta)$ $=\nabla^{2}\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.\vspace{-1mm}$
Then, the design of the OMD algorithm can be conceptually divided into two parts: the approximation of the past losses and the approximation of the current loss. We provide the details of each below.
Approximate the past losses.  To integrate historical information from previous iterations while avoiding the use of MLE in Eq. (1), we construct the estimator $\bar{\theta}_{k+1,h}$ using the implicit OMD form:
$\bar{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\ell_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\bar{\theta}_{k,h}\rVert_{\bar{\mathcal{H}}_{k,h}}^{2}\Big{\}},$ (8)
where $\eta$ is a step size and ¯$\bar{\mathcal{H}}_{k,h}\triangleq\bar{\mathcal{H}}_{k,h}(\bar{\theta}_{k+1,h})=\sum_{i=1}^{k-1}H_{i,h}(\bar{\theta}_{i+1,h})+\lambda_{k}I$. The optimization problem can be decomposed in two terms. The first term is the instantaneous log-loss $\ell_{k,h}(\theta)$, which accounts for the information of the current episode. The second is a regularization term that ensures the current model remains close to the previous one, $\bar{\theta}_{k,h}$, thereby incorporating the historical information acquired so far. The most critical aspect in the above is the design of the local norm ¯$\bar{\mathcal{H}}_{k,h}$, which intentionally approximate the per-episode Hessian matrix by $H_{i}(\bar{\theta}_{i+1,h})$ at a look ahead point $\bar{\theta}_{i+1,h}$. Such a Hessian matrix, originally introduced by Faury et al. (2022), effectively captures the local curvature of the loss function and is crucial for ensuring statistical efficiency.
The update rule in Eq. (8) is storage efficient, as it only requires storing the Hessian matrix ¯$\bar{\mathcal{H}}_{k,h}$, which can be updated incrementally, resulting in an $\mathcal{O}(1)$ storage cost. In terms of time complexity, the optimization problem in Eq. (8) suffers an $\mathcal{O}(\log k)$ time complexity at episode $k$, since the loss function is defined only over the current episode. While this represents a significant improvement over the $\mathcal{O}(k\log k)$ time complexity of the MLE in Eq. (1), there is still a need to reduce the cost further to $\mathcal{O}(1)$ per episode, particularly given the potentially large number of episodes.
Approximate the current loss.  To achieve $\mathcal{O}(1)$ time complexity per episode, we can further approximate the current loss with a second order approximation. Drawing inspiration from Zhang and Sugiyama (2023), we define the second-order approximation of the original loss function $\ell_{k,h}(\theta)$ at $\widetilde{\theta}_{k,h}$ as ~$\widetilde{\ell}_{k,h}(\theta)=\ell_{k,h}(\widetilde{\theta}_{k,h})+\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})}^{2}$, where $\widetilde{\theta}_{k,h}$ is the current estimate. Then, we can replace $\ell_{k,h}(\theta)$ with its second-order approximation ~$\widetilde{\ell}_{k,h}(\theta)$ in the optimization problem in Eq. (8). This leads to the following approximate optimization problem:
@@ 5.1 Efficient Online Parameter Estimation
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}\Big{\}}.$ (9)
where $\eta$ is the step size, ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$ and $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda_{k}I$. Then, Eq. (9) can be solved with a single projected gradient step with the following equivalent formulation:
$\widetilde{\theta}^{\prime}_{k+1,h}=\widetilde{\theta}_{k,h}-\eta\widetilde{\mathcal{H}}_{k,h}^{-1}\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}.$
Thus, Eq. (9) is computationally efficient, as it only suffers an $\mathcal{O}(1)$ storage and time complexity.
Notice that the update rule in Eq. (9) is actually a standard online mirror descent (OMD) formula,
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta\rangle+\frac{1}{\eta}\mathcal{D}_{\psi_{k}}(\theta,\widetilde{\theta}_{k,h})\Big{\}}.$ (10)
where the regularizer is $\psi_{k}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}$ and $\mathcal{D}_{\psi_{k}}(\cdot,\cdot)$ is the induced Bregman divergence. Therefore, we can construct the confidence set building upon the modern analysis of OMD (Orabona, 2019; Zhao et al., 2024). Specifically, we can construct the $\kappa$-independent confidence set as follows.
@@
For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, define
@@
Compared to the confidence set in Lemma 1, the radius $\widetilde{\beta}_{k}$ in Lemma 2 includes an additional $\log U$ factor. This is due to our approximation of the original MLE using the OMD update.
5.2 Efficient Optimistic Value Function Construction
Although the optimistic value function in Eq. (6) preserves local information effectively and provides strong theoretical guarantees, it is computationally intractable due to the need to solve a non-convex optimization problem. To address this challenge, we propose an efficient method in this section.
The key idea is to use a second-order Taylor expansion to derive a closed-form bonus term, which replaces the operation of taking the maximum over the non-convex confidence set. While this idea has been used in bandit settings, fundamental challenges arise when applying it in the MDP setting. Specifically, Zhang and Sugiyama (2023) studied the multi-parameter MLogB bandit, where each outcome is associated with a distinct parameter vector. In contrast, MNL mixture MDPs involve a single shared parameter vector across all outcomes. This distinction leads to a more complex Hessian matrix, necessitating a more sophisticated analysis. A direct use of their analysis will leads to a polynomial dependence on the number of reachable states $U$, which is undesirable in the MDP setting. Lee and Oh (2024) focused on the single-parameter MNL bandit, which is more closely related to our setting. However, they construct the optimistic value function by directly taking the maximum over the confidence set, a computationally intractable approach in the MDP setting. As a result, they can apply a second-order Taylor expansion around the ground truth parameter $\theta_{h}^{*}$ in their analysis, while we must apply it around the estimated parameter $\widetilde{\theta}_{k,h}$ to construct the bonus term explicitly.
Lemma 3.
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}.$
where
$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]},$ (11)
where $\widetilde{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s,a)$. In contrast to the value function in Eq. (5), which incorporates the term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$, the refined value function in Eq. (11) replaces it with $\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}$. This modification better preserves local information, offering a more accurate estimation error bound.
5.3 Regret Guarantee
Algorithm 2 UCRL-MNL-OL
@@ 5.3 Regret Guarantee
3:     Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in Eq. (11).
@@
7:        Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}$.
8:        Update $\mathcal{H}_{k+1,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
@@
The overall algorithm UCRL-MNL-OL is similar to UCRL-MNL-LL, but with the estimator and optimistic value function updated in a computationally efficient manner. The detailed algorithm is presented in Algorithm 2. We provide the guarantee of UCRL-MNL-OL in the following theorem.
@@
For any $\delta\in(0,1)$, set $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(kH/\delta))$, $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, with probability at least $1-\delta$, UCRL-MNL-LL (Algorithm 2) ensures
@@
UCRL-MNL-OL attains the same regret as UCRL-MNL-LL, but with constant computational cost per episode. This is achieved by constructing an efficient online estimation based on OMD and an optimistic value function by closed-form bonus instead of the non-convex optimization.